# Phenome-Wide Multi-Omics Integration Uncovers Distinct Archetypes of Human Aging

### v2

@@ Abstract
Aging is a highly complex and heterogeneous process that progresses at different rates across individuals, making biological age (BA) a more accurate indicator of physiological decline than chronological age. While previous studies have built aging clocks using single-omics data, they often fail to capture the full molecular complexity of human aging. In this work, we leveraged the Human Phenotype Project [^1] https://knowledgebase.pheno.ai/, a large-scale cohort of 12,000 adults aged 30–70 years, with extensive longitudinal profiling that includes clinical, behavioral, environmental, and multi-omics datasets—spanning transcriptomics, lipidomics, metabolomics, and the microbiome. By employing advanced machine learning frameworks capable of modeling nonlinear biological dynamics, we developed and rigorously validated a multi-omics aging clock that robustly predicts diverse health outcomes and future disease risk. Unsupervised clustering of the integrated molecular profiles from multi-omics uncovered distinct biological subtypes of aging, revealing striking heterogeneity in aging trajectories and pinpointing pathway-specific alterations associated with different aging patterns. These findings demonstrate the power of multi-omics integration to decode the molecular landscape of aging and lay the groundwork for personalized healthspan monitoring and precision strategies to prevent age-related diseases.
@@ 2.1 System-wide characterization in a large healthy aging cohort
Figure 1: Comprehensive data collected in an aging cohort. a, Illustration of the comprehensive clinical, physiological and behavioral data collected in the 10K study which are assigned into 12 system-level categories. b, Sample collection and multi-omics data acquisition of the cohort. Three types of biological samples were collected, and 5 types of omics data were acquired. c, Spearman correlation between the first principal component and ages for each type of systems. d, Heatmap depicting the dynamic changing molecules and microbes during human aging.
To explore the phenotypes of various systems during aging, we analyzed an ongoing collection project HPP, which is a large-scale phenotyped cohort consisting of ethnically diverse people in Israel, most of whom are healthy and have received higher education. This deep phenotyped cohort includes more than 12,000 participants aged 30-70 years with a 5-year follow-up and consist of not only multi-omics data but also clinical, physiological, and behavioral data, which can be divided into 12 body systems (Fig. 1a). The BMI of the participants ranged from 15.73 to 47.23. Among the participants, 47.49% were male. For each follow-up, the blood, stool, and oral swab samples were collected. A total of 31,433 biological samples (including 11,923 blood samples, 10,694 stool samples, and 8,816 oral swab samples) were collected. The biological samples were used for multi-omics profiling (including transcriptome from peripheral blood mononuclear cells, metabolomics in plasma, lipidomics in plasma, gut microbiome, and oral microbiome). A total of 52,874 biological features (including 49,614 transcripts, 250 metabolites, 187 lipids, 736 gut microbiome taxa, and 2,087 oral microbiome taxa) were collected, generating 9,787,326,308,545 data points.
To elucidate the multi-system effects of aging, we examined the association between HPP data and physiological age. Most systemic indicators displayed significant age dependency (Fig. 1.c), with cardiovascular and sleep measures showing the strongest positive correlations with age. These findings reinforce the view that aging is a complex, multisystem process and underscore the utility of our HPP dataset for quantifying and tracking human aging trajectories. Moreover, biological features exhibited pronounced nonlinear changes across age (Fig. 1d). Together, this extensive multi-omics resource enables comprehensive investigation of molecular, phenotypic, and microbial dynamics during aging. The large cohort (12,000 participants) provides exceptional power to detect intricate aging-related changes across both molecular and functional dimensions.
@@ 2.2 Multi-omics aging clocks predict biological age
Figure 2: Internal validation demonstrates the accuracy, robustness, and systemic relevance of multi-omics aging clocks. a, Pearson correlation (two-sided test) is plotted between actual chronological age and predicted age for each omics type. Linear fit is plotted as a dashed line. b, To verify the robustness of the aging clock models, the training set and internal validation set are randomly resampled, and the model is trained for 10 times. Pearson correlation (two-sided test), R2, r.m.s.e., m.a.e. and med.a.e are used as evaluation criteria. c, Proportion of biomarkers significantly associated with each omics AgeAccel across different physiological systems. For each system, elements represent the percentage of all measured biomarkers within that system that are significantly correlated (FDR-adjusted p << 0.05, Pearson correlation) with a given omics AgeAccel. This analysis highlights the systemic breadth of each clock. d, UpSet plot showing the number of unique and shared biomarkers significantly associated with the different aging omics AgeAccels. The horizontal bars indicate the total number of significant biomarkers for each omics AgeAccel. The upper vertical bars show the number of biomarkers in each intersection, with the matrix below indicating which omics AgeAccels are included in that intersection. This illustrates the specificity and overlap of biological signals captured by each type of omics AgeAccels.
@@
We developed a multi-omics aging clock using five types of omics data derived from human PBMC, feces, and saliva samples. During the training phase, we compared three machine learning methods (least absolute shrinkage selection operator (LASSO) [26], Elastic Net [27], LightGBM [28]) to train multi-omics age clock models. We evaluated the performance of the aging clock by comparing the predicted age with the CA (Pearson’s R, R2, root mean square error (r.m.s.e.), median absolute error (m.a.e.) and median absolute error (med.a.e)). A cross-validation step was implemented, where 80% of the samples were randomly selected for model training, while the remaining 20% of the data were used as an internal validation set. Fig. 2a underscored the accuracy of our aging clocks. In our initial experiments, we found that LightGBM showed superior performance across all omics types. Consequently, we selected it as the final model and used SHapley Additive exPlanations (SHAP) [29] to identify omics features associated with predicting chronological age. The strong correlation between chronological age and predicted transcriptional age demonstrated the accuracy of our aging clock. We further confirmed the robustness of the aging clock model through 10-fold cross-validation, which showed only slight performance changes in all metrics (Fig. 2b).
@@
To evaluate the physiological relevance of age acceleration across biological strata, we assessed associations between five omics-based age-acceleration (AgeAccel) scores and a comprehensive panel of parameters spanning 12 physiological and lifestyle domains (Fig. 2c). AgeAccel scores, computed as residuals from regressing biological age on chronological age, differed substantially in their links to systemic health, after adjusting for multiple hypotheses (false discovery rate (FDR) << 0.05). Of these, lifestyle was significantly associated with all five omics-based AgeAccel metrics (propprop=0.10 for transcriptomics, propprop=0.01 for metabolomics, propprop=0.28 for lipidomics, propprop=0.78 for gut microbiome and propprop=0.77 for oral microbiome). Gut-microbiome AgeAccel showed the broadest association profile (significant associations with Liver health, Insulin resistance, frailty, lifestyle and hematopoietic system) and was the only omics-derived AgeAccel associated with frailty,onsistent with previous findings by [30]. Oral-microbiome AgeAccel showed a similar but weaker pattern. The oral-microbiome AgeAccel showed the strongest association with DXA-derived bone density (propprop=0.24), independent of sex and BMI. This pattern aligns with previous reported links between periodontal dysbiosis and systemic bone loss [31, 32, 33]. Moreover, oral microbiome composition significantly shifts (P << 2.2e-16) with age and reflects aspects of biological aging, providing a plausible rationale for an AgeAccel–bone density association, which is aligned with [34]. In contrast to the broad profiles of the microbiome-based metrics, the remaining molecular AgeAccel clocks exhibited more circumscribed physiological associations. The Transcriptomic-AgeAccel was predominantly linked to metabolic dysregulation, with biomarkers of insulin resistance constituting nearly half of its significant associations (propprop=0.45). Weaker, yet significant, links were also observed with diet (propprop=0.11), sleep (propprop=0.11), and lifestyle (propprop=0.10), while its connection to the cardiovascular system was minimal (propprop=0.02). The Metabolomic-AgeAccel displayed an even more specific signature, being almost exclusively associated with the renal function system (propprop=0.50). Conversely, the Lipidomic-AgeAccel captured a more pleiotropic signal; it was most strongly dominated by associations with diet (propprop=0.67) but was also substantially linked to the hematopoietic (propprop=0.43) and immune (propprop=0.40) systems, in addition to lifestyle (propprop=0.28).
@@
We quantified specificity and redundancy across AgeAccel metrics by analyzing set intersections of FDR-significant associations (Fig. 2.d). After controlling for multiple testing, each AgeAccel retained a non-overlapping subset of parameters, with the oral-microbiome AgeAccel showing the largest unique subset (108 of 251 parameters), consistent with prior evidence that the oral microbiome captures aging-related and systemic signals [35, 36]. Beyond unique signals, we detected non-random overlaps: the largest intersection involved parameters associated with both gut and oral-microbiome AgeAccel, in line with microbiome-based aging clocks and reports on the oral–gut axis [37, 38]. Furthermore, a smaller but crucial set of biomarkers was significantly associated with four aging AgeAccels, suggesting a conserved, multi-system aging signal that is detectable across most biological layers of organization.
@@ 2.5 Multi-omics aging clocks predict multimorbidity
To further understand the impact of age-related chronic inflammation on age-related pathologies, we calculated regression analyses between the total number of age-related diseases (multimorbidity) and the multi-omics aging clock. Multimorbidity is a global health research priority and has become the gold standard for aging research because it represents the accumulation of physiological damage in an individual[45, 46, 47, 48, 49]. All of these disease features were binary. For these analyses, we controlled for sex, BMI, as each of these variables has reported effects in the etiology of age-related pathologies. We observed a significant correlation between the multi-omics aging clock and multimorbidity in the elderly in this study (>> 60 years old) (Fig. 5a). This highlights the key role of the multi-omics aging clock in identifying the accumulation of physiological damage during aging.
@@
Five different omics AgeAccel scores and the incidence of 16 common non-cancer diseases showed different specificities. Since our study has not yet been followed up for a long time, we examined whether the calculated AgeAccel scores were associated with medical diagnosis at baseline. We used a logistic regression model to estimate the association between the five AgeAccel scores and 16 common non-cancer diseases after adjusting for actual age, sex, and BMI. The analysis results revealed that the association between the five omics aging AgeAccel scores and diseases showed high specificity, among which the microbiome AgeAccel scores showed the most extensive and strongest association (Fig. 5b). The gut microbiome AgeAccel score showed the strongest disease association signal among all AgeAccel scores, and was statistically significantly positively correlated with 14 of the 16 diseases. The strongest association was in metabolic diseases. Only fractures and migraines did not show significant associations (Fig. 5b). Similarly, the oral microbiome AgeAccel score also showed a wide range of associations, significantly associated with 10 of the 16 diseases. Unlike the gut microbiome, its association with osteoarthritis, hemorrhoids, migraine, hypothyroidism, headaches and fractures is not statistically significant (Fig. 5b). Compared with the microbiome AgeAccel score, the molecular omics AgeAccel score derived from blood has a narrower range of associations, mainly concentrated in cardiometabolic diseases. The transcriptome AgeAccel score has the weakest association signal, with only a weak but statistically significant association with hypothyroidism and hyperglycemia (Fig. 5b).
@@ 3 Discussion
A key innovation of our work is the characterization of distinct aging types, accelerated and decelerated aging—based on cross-sectional multi-omics data (Fig. 3a,b,c,g). We demonstrate that accelerated aging is not simply a faster version of normative aging but a distinct process marked by an early and large-scale disruption of molecular homeostasis, with the most pronounced dysregulation originating in the microbiome. This disruption manifests as precipitous changes in molecular trajectories during mid-life, typically beginning around age 50. In contrast, decelerated aging is defined by prolonged molecular stability and more gradual, later-life changes. This distinction suggests that a critical window for interventions aimed at promoting healthy aging may lie in mid-life, prior to this sharp inflection point.
@@
From a clinical standpoint, our multi-omics aging clocks demonstrate clear relevance to human health and disease (Fig. 5a,b). The biological age scores, particularly from the gut microbiome clock, were significantly associated with multimorbidity in older adults and showed strong, specific associations with the risk of 14 out of 16 common non-cancer diseases, most notably metabolic disorders such as obesity, fatty liver, and hypertension. These associations remained robust after adjusting for chronological age, sex, and BMI, indicating that our clocks capture a dimension of physiological dysregulation and accumulated damage not reflected by traditional risk factors. This positions the multi-omics biological age as a promising surrogate endpoint for clinical trials and a powerful tool for personalized risk stratification and preventive medicine.
@@ 4.1 Description of cohort
The data for this study were derived from a prospective longitudinal cohort of the Human Phenotype Project (HPP), which was launched in Israel in 2018. We included participants from this cohort who completed the baseline visit between January 2019 and December 2022. After data quality control, a total of 10,019 participants were included in the construction and validation analysis of this aging clock. These participants ranged in age from 30 to 70 years old, provided complete clinical phenotype data at the baseline visit, and successfully completed the collection of biological samples for multi-omics analysis. This study strictly adhered to the ethical principles of the Declaration of Helsinki and was approved by the Institutional Review Board of the Weizmann Institute of Science in Israel (IRB approval number: 1719-1). All participants were fully informed of the study content and signed informed consent before joining the study. This cohort study is designed to track participants for up to 25 years, with systematic follow-up planned every 2 years. At recruitment, individuals with severe or chronic illnesses were excluded from the study, providing an ideal baseline for studying the normative aging process in a relatively healthy population. The longitudinal nature of this design, with a healthy baseline population, is critical for developing biological age clocks that can capture differences in aging rates and predict future health risks.

### v3

@@ Abstract
Aging is a highly complex and heterogeneous process that progresses at different rates across individuals, making biological age (BA) a more accurate indicator of physiological decline than chronological age. While previous studies have built aging clocks using single-omics data, they often fail to capture the full molecular complexity of human aging. In this work, we leveraged the Human Phenotype Project [^1] https://knowledgebase.pheno.ai/, a large-scale cohort of 10,000 adults aged 40–70 years, with extensive longitudinal profiling that includes clinical, behavioral, environmental, and multi-omics datasets—spanning transcriptomics, lipidomics, metabolomics, and the microbiome. By employing advanced machine learning frameworks capable of modeling nonlinear biological dynamics, we developed and rigorously validated a multi-omics aging clock that robustly predicts diverse health outcomes and future disease risk. Unsupervised clustering of the integrated molecular profiles from multi-omics uncovered distinct biological subtypes of aging, revealing striking heterogeneity in aging trajectories and pinpointing pathway-specific alterations associated with different aging patterns. These findings demonstrate the power of multi-omics integration to decode the molecular landscape of aging and lay the groundwork for personalized healthspan monitoring and precision strategies to prevent age-related diseases.
@@ 1 Introduction
Figure 1: Comprehensive data collected in an aging cohort. a, Illustration of the comprehensive clinical, physiological and behavioral data collected in the 10K study which are assigned into 12 system-level categories. b, Collection time range and sample numbers for the cohort. c, Sample collection and multi-omics data acquisition of the cohort. Three types of biological samples were collected, and 5 types of omics data were acquired. d, Spearman correlation between the first principal component and chronological ages for each type of systems. e, Heatmap depicting the dynamic changing molecules and microbes during human aging.
@@ 2.1 System-wide characterization in a large healthy aging cohort
To explore the phenotypes of various systems during aging, we analyzed an ongoing collection project HPP, which is a large-scale phenotyped cohort consisting of ethnically diverse people in Israel, most of whom are healthy and have received higher education. This deep phenotyped cohort includes more than 10,000 participants aged 40-70 years with a 5-year follow-up and consist of not only multi-omics data but also clinical, physiological, and behavioral data, which can be divided into 12 body systems (Fig. 1a, b). The BMI of the participants ranged from 15.73 to 47.23. Among the participants, 47.49% were male. For each follow-up, the blood, stool, and oral swab samples were collected. A total of 31,433 biological samples (including 11,923 blood samples, 10,694 stool samples, and 8,816 oral swab samples) were collected (Fig. 1c). The biological samples were used for multi-omics profiling (including transcriptome from peripheral blood mononuclear cells, metabolomics in plasma, lipidomics in plasma, gut microbiome, and oral microbiome). A total of 52,874 biological features (including 49,614 transcripts, 250 metabolites, 187 lipids, 736 gut microbiome taxa, and 2,087 oral microbiome taxa) were collected, generating 9,787,326,308,545 data points.
To elucidate the multi-system effects of aging, we examined the association between HPP data and physiological age. Most systemic indicators displayed significant age dependency (Fig. 1d), with cardiovascular and sleep measures showing the strongest positive correlations with age. These findings reinforce the view that aging is a complex, multisystem process and underscore the utility of our HPP dataset for quantifying and tracking human aging trajectories. Moreover, biological features exhibited pronounced nonlinear changes across age (Fig. 1e). Together, this extensive multi-omics resource enables comprehensive investigation of molecular, phenotypic, and microbial dynamics during aging. The large cohort (12,000 participants) provides exceptional power to detect intricate aging-related changes across both molecular and functional dimensions.
@@ 2.2 Multi-omics aging clocks predict biological age
Figure 2: Internal validation demonstrates the accuracy, robustness, and systemic relevance of multi-omics aging clocks. a, Pearson correlation (two-sided test) is plotted between actual chronological age and predicted age for each omics type. Linear fit is plotted as a dashed line. b, To verify the robustness of the aging clock models, the training set and internal validation set are randomly resampled, and the model is trained for 10 times. Pearson correlation (two-sided test), $r^{2}$, r.m.s.e., m.a.e. and med.a.e are used as evaluation criteria. c, Proportion of biomarkers significantly associated with each omics AgeAccel across different physiological systems. For each system, elements represent the percentage of all measured biomarkers within that system that are significantly correlated (FDR-adjusted p << 0.05, Pearson correlation) with a given omics AgeAccel. This analysis highlights the systemic breadth of each clock. d, UpSet plot showing the number of unique and shared biomarkers significantly associated with the different aging omics AgeAccels. The horizontal bars indicate the total number of significant biomarkers for each omics AgeAccel. The upper vertical bars show the number of biomarkers in each intersection, with the matrix below indicating which omics AgeAccels are included in that intersection. This illustrates the specificity and overlap of biological signals captured by each type of omics AgeAccels.
@@
We developed a multi-omics aging clock using five types of omics data derived from human PBMC, feces, and saliva samples. During the training phase, we compared three machine learning methods (least absolute shrinkage selection operator (LASSO) [26], Elastic Net [27], LightGBM [28]) to train multi-omics age clock models. We evaluated the performance of the aging clock by comparing the predicted age with the CA (Pearson’s R, $r^{2}$, root mean square error (r.m.s.e.), median absolute error (m.a.e.) and median absolute error (med.a.e)). A cross-validation step was implemented, where 80% of the samples were randomly selected for model training, while the remaining 20% of the data were used as an internal validation set. Fig. 2a underscored the accuracy of our aging clocks. In our initial experiments, we found that LightGBM showed superior performance across all omics types. Consequently, we selected it as the final model and used SHapley Additive exPlanations (SHAP) [29] to identify omics features associated with predicting chronological age. The strong correlation between chronological age and predicted transcriptional age demonstrated the accuracy of our aging clock. We further confirmed the robustness of the aging clock model through 10-fold cross-validation, which showed only slight performance changes in all metrics (Fig. 2b).
@@
To evaluate the physiological relevance of age acceleration across biological strata, we assessed associations between five omics-based age-acceleration (AgeAccel) scores and a comprehensive panel of parameters spanning 12 physiological and lifestyle domains (Fig. 2c). AgeAccel scores, computed as residuals from regressing biological age on chronological age, differed substantially in their links to systemic health, after adjusting for multiple hypotheses (false discovery rate (FDR) << 0.05). Of these, diet, lifestyle and hematopoietic system were significantly associated with four omics-based AgeAccel metrics. Gut-microbiome AgeAccel showed the broadest association profile (significant associations with diet, liver health, insulin resistance, frailty, lifestyle and hematopoietic system) and was the only omics-derived AgeAccel associated with frailty, consistent with previous findings by [30]. Oral-microbiome AgeAccel showed a similar but weaker pattern. The oral-microbiome AgeAccel showed the strongest association with diet (propprop=0.78), independent of sex and BMI. This pattern aligns with previous reported links whereby diet shapes the oral microbiota by providing nutrients and selecting organisms adapted to specific dietary resources [31, 32]. In contrast to the broad profiles of the microbiome-based metrics, the remaining molecular AgeAccel clocks exhibited more circumscribed physiological associations. The Transcriptomic-AgeAccel was predominantly linked to immune and inflammatory processes, with biomarkers of the hematopoietic system accounting for about one-third of its significant associations. Weaker, yet significant, links were also observed with diet (propprop=0.11), insulin resistance (propprop=0.15), and lifestyle (propprop=0.06), while its connection to the cardiovascular system was minimal (propprop=0.02). The Metabolomic-AgeAccel displayed an even more specific signature, being almost exclusively associated with the hematopoietic system (propprop=0.43). Conversely, the lipidomic-AgeAccel captured a more pleiotropic signal; it was most strongly dominated by associations with diet (propprop=0.89) but was also substantially linked to the lifestyle (propprop=0.33) and immune (propprop=0.30) systems in addition to hematopoietic system (propprop=0.57).
@@
We quantified specificity and redundancy across AgeAccel metrics by analyzing set intersections of FDR-significant associations (Fig. 2d). After controlling for multiple testing, each AgeAccel retained a non-overlapping subset of parameters, with the gut-microbiome AgeAccel showing the largest unique subset (104 of 227 parameters), consistent with prior evidence that the gut microbiome captures aging-related and systemic signals [33, 34]. Beyond unique signals, we detected non-random overlaps: the largest intersection involved parameters associated with both gut and oral-microbiome AgeAccel, in line with microbiome-based aging clocks and reports on the oral–gut axis [35, 36]. Furthermore, a smaller but crucial set of biomarkers was significantly associated with four aging AgeAccels, suggesting a conserved, multi-system aging signal that is detectable across most biological layers of organization.
@@ 2.5 Multi-omics aging clocks predict multimorbidity
To further understand the impact of age-related chronic inflammation on age-related pathologies, we calculated regression analyses between the total number of age-related diseases (multimorbidity) and the multi-omics aging clock. Multimorbidity is a global health research priority and has become the gold standard for aging research because it represents the accumulation of physiological damage in an individual[43, 44, 45, 46, 47]. All of these disease features were binary. For these analyses, we controlled for sex, BMI, as each of these variables has reported effects in the etiology of age-related pathologies. We observed a significant correlation between the multi-omics aging clock and multimorbidity in the elderly in this study (Fig. 5a). This highlights the key role of the multi-omics aging clock in identifying the accumulation of physiological damage during aging.
@@
Five different omics AgeAccel scores and the incidence of 16 common non-cancer diseases showed different specificities. Since our study has not yet been followed up for a long time, we examined whether the calculated AgeAccel scores were associated with medical diagnosis at baseline. We used a logistic regression model to estimate the association between the five AgeAccel scores and 16 common non-cancer diseases after adjusting for actual age, sex, and BMI. The analysis results revealed that the association between the five omics aging AgeAccel scores and diseases showed high specificity, among which the microbiome AgeAccel scores showed the most extensive and strongest association (Fig. 5b). The gut microbiome AgeAccel score showed the strongest disease association signal among all AgeAccel scores, and was statistically significantly positively correlated with 13 of the 16 diseases. The strongest association was in obesity diseases. Only asthma, cholelithiasis and fractures did not show significant associations (Fig. 5b). The oral microbiome AgeAccel score also showed a similar but weaker range of associations, significantly associated with 7 of the 16 diseases. Unlike the gut microbiome, its association with cholelithiasis is statistically significant (Fig. 5b). Compared with the microbiome AgeAccel score, the molecular omics AgeAccel score derived from blood has a narrower range of associations, mainly concentrated in cardiometabolic diseases. The transcriptome AgeAccel score has the weakest association signal, with only a weak but statistically significant association with hyperglycemia (Fig. 5b).
@@ 3 Discussion
A key innovation of our work is the characterization of distinct aging types, accelerated and decelerated aging—based on cross-sectional multi-omics data (Fig. 3a,b,c,d). We demonstrate that accelerated aging is not simply a faster version of normative aging but a distinct process marked by an early and large-scale disruption of molecular homeostasis, with the most pronounced dysregulation originating in the microbiome. This disruption manifests as precipitous changes in molecular trajectories during mid-life, typically beginning around age 50. In contrast, decelerated aging is defined by prolonged molecular stability and more gradual, later-life changes. This distinction suggests that a critical window for interventions aimed at promoting healthy aging may lie in mid-life, prior to this sharp inflection point.
@@
From a clinical standpoint, our multi-omics aging clocks demonstrate clear relevance to human health and disease (Fig. 5a,b). The biological age scores, particularly from the gut microbiome clock, were significantly associated with multimorbidity in older adults and showed strong, specific associations with the risk of 13 out of 16 common non-cancer diseases, most notably metabolic disorders such as obesity, fatty liver, and hypertension. These associations remained robust after adjusting for chronological age, sex, and BMI, indicating that our clocks capture a dimension of physiological dysregulation and accumulated damage not reflected by traditional risk factors. This positions the multi-omics biological age as a promising surrogate endpoint for clinical trials and a powerful tool for personalized risk stratification and preventive medicine.
@@ 4.1 Description of cohort
The data for this study were derived from a prospective longitudinal cohort of the Human Phenotype Project (HPP), which was launched in Israel in 2018. We included participants from this cohort who completed the baseline visit between February 2019 and January 2023. After data quality control, a total of 10,019 participants were included in the construction and validation analysis of this aging clock. These participants ranged in age from 40 to 70 years old, provided complete clinical phenotype data at the baseline visit, and successfully completed the collection of biological samples for multi-omics analysis. This study strictly adhered to the ethical principles of the Declaration of Helsinki and was approved by the Institutional Review Board of the Weizmann Institute of Science in Israel (IRB approval number: 1719-1). All participants were fully informed of the study content and signed informed consent before joining the study. This cohort study is designed to track participants for up to 25 years, with systematic follow-up planned every 2 years. At recruitment, individuals with severe or chronic illnesses were excluded from the study, providing an ideal baseline for studying the normative aging process in a relatively healthy population. The longitudinal nature of this design, with a healthy baseline population, is critical for developing biological age clocks that can capture differences in aging rates and predict future health risks.